OpenStudy (anonymous):

Put the equation in slope-intercept form and identify the slope of the line. 3x + 9y = 12

OpenStudy (uri):

y=mx+b arrange it in this form

OpenStudy (uri):

9y=-3x+12 y=-3/9x +12/9

OpenStudy (texaschic101):

slope intercept form : y = mx + b step 1 -- subtract 3x from both sides step 2 -- divide both sides by 9 step 3 -- reduce the slope will be in the m position

OpenStudy (anonymous):

See, that's what I don't understand. My teacher has not explained what exactly it means to do something "on both sides."

OpenStudy (uri):

Subract 3 from both sides means 3x + 9y = 12 -3 -3

OpenStudy (anonymous):

So it's 3x - 3 and 12 - 3